OpenStudy (anonymous):

What is the value of x 5x + 2 = x + 12 with working out

OpenStudy (anonymous):

5x-x=12-2 4x=10 x=5/2=2.5

OpenStudy (angela210793):

5x+2-x-12=0 4x-10=0 4x=10 x=10/4=5/2

OpenStudy (anonymous):

alright thank you :D

OpenStudy (angela210793):

Uw :)

OpenStudy (anonymous):

Can anyone work out the value of y in this? 4 - 2y = y + 2

OpenStudy (angela210793):

4 - 2y = y + 2 4-2=y+2y 2=3y y=3/2

OpenStudy (aroub):

4-2y-y-2=0 2-3y=0 -3y=-2 y=2/3

OpenStudy (angela210793):

2/3 right....sorry :)

OpenStudy (aroub):

:)

OpenStudy (anonymous):

\[2/3\]

OpenStudy (anonymous):

i got 1.5 somehow :S

OpenStudy (angela210793):

u must have got 3/2 as i did b4...but it's 2/3 and u got the answer...check it